# RAUCA: A Novel Physical Adversarial Attack on Vehicle Detectors via Robust and Accurate Camouflage Generation

### v1

@@ 5 Conclusion
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Datasets
In this section we present the parameters of all of the datasets in our experiments.
@@
For NRP model training and testing: We collect vehicle datasets in the CARLA simulation environment under 16 distinct weather conditions. These conditions are created by combining four sun altitude angles (-90, -30, 30, 90) with four fog densities (0, 30, 60, 90). Within each weather scenario, we randomly choose 30 locations for model training and 3 locations for testing. For each car locations, imagery is acquired at two azimuth angles (0, 45 degrees), two altitude angles (22.5, 67.5 degrees), and four distances (5, 10, 15, 20 meters) . In terms of color paint for the vehicle, nine vehicle colors are utilized during training(red, green, blue, magenta, yellow, cyan, white, and black and grey ) and 64 vehicle colours (RGB values taken from [0, 85, 170, 255] respectively) for testing. In summary, we employ 69,120 photo-realistic images for model training (consisting 16 weather conditions $\times$ 20 car locations $\times$ 16 camera poses $\times$ 9 colors) and 49152 for testing (consisting 16 weather conditions $\times$ 3 car locations $\times$ 16 camera poses $\times$ 64 colors)
@@
For texture generation and evaluation in physically-based simulation settings: For the generation of adversarial camouflage, we choose the same 16 weather combinations as NRP training. For evaluation in physically-based simulation, we produce two datasets, one with the same weather combinations as the training set, the other is featuring 16 novel weather conditions that are not seen during the texture generation, including diverse scenarios like MidRainyNight and DustStorm. We use the former dataset for robust evaluation and the latter one for transferability evaluation. of our adversarial camouflage. Within each weather scenario, we randomly choose 20 locations for texture generation and 4 distinct locations for each test dataset. For each car transformation, imagery is captured at every 45-degree increment in azimuth angle, spanning four distinct altitude angles (0.0, 22.5, 45.0, 67.5 degrees) and four varying distances (5, 10, 15, 20 meters) for each. In summary, we employ 40,960 photo-realistic images for model training (consisting 16 weather conditions $\times$ 20 car locations $\times$ 128 camera poses $\times$ 1 colors) and 8192 for each test dataset (consisting 16 weather conditions $\times$ 4 car locations $\times$ 128 camera poses $\times$ 1 colors).
@@
For evaluation in the real world settings: We print our textures and adversarial textures from DAS (Wang et al., 2021), FCA (Wang et al., 2022), DTA (Suryanto et al., 2022) as well as ACTIVE (Suryanto et al., 2023) and stick them on official 1:12 Audi E-Tron vehicle models prepare an unpainted. In attition to this, We also prepare an unpainted vehicle. We shoot these models in the real world with the Xiaomi 14 for comparison. To demonstrate the robustness of our camouflage in physical muli-light conditions, we capture photos at different times to change the light condition. We take 2592 pictures of six cars with poses of 3 distances (5cm, 10cm, 15cm), 8 azimuths per 45°, and 3 altitude angles (0,30,60), as well as at two locations at noon, in the afternoon and at night . Moreover, we evaluate textures effectiveness in the presence of fog in the real world. We use a smoke machine to simulate the foggy day. We capture pictures of six cars at every 45-degree increment in azimuth angle, 3 altitude angles (0, 30, 60 degrees). We take five consecutive pictures for the same camera angle for each car to reduce the unfairness caused by the randomness of the fog. Finally, we obtained 720 pictures for comparison.

### v2

@@ 5 Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grants: 62306093, 62376074), the National Key R&D Program of China (Grant: 2021YFB2700900), the Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies (Grant: 2022B1212010005), and Shenzhen Science and Technology Program (Grants: JSGGKQTD20221101115655027, KCXST20221021111404010, JSGG20220831103400002, SGDX20230116091244004, KJZD20230923114405011, KJZD20231023095959002, RKX20231110090859012), the Fundamental Research Funds for the Central Universities (Grants: HIT.DZJJ.2023118, HIT.OCEF.2024047), and the Fok Ying Tung Education Foundation of China (Grant: 171058).
Impact Statement
Attackers might exploit our proposed adversarial camouflage method to attack autonomous driving and transportation systems. To remedy this, we suggest enhancing the current vehicle detection models via adversarial training with our camouflage, which can reinforce models against these potential attacks. Additionally, the generated adversarial camouflage can be used to test the models and ensure that the model is robust before deployment.
@@ Appendix A Datasets
This section presents the parameters of all of the datasets in our experiments.
@@
For NRP model training and testing: We collect vehicle datasets in the CARLA simulation environment under 16 weather conditions. These weather conditions are generated by combining four sun altitude angles (-90°, -30°, 30°, 90°) with four fog densities (0, 25, 50, 90). Within each weather scenario, we randomly choose 30 locations for model training and 3 for testing. For each car location, imagery is acquired at two azimuth angles (0°, 45°), two altitude angles (22.5°, 67.5°), and four distances (5m, 10m, 15m, 20m). In terms of color paint for the vehicle, nine vehicle colors are utilized during training(red, green, blue, magenta, yellow, cyan, white, and black and grey ), and 64 vehicle colors (RGB values are taken from [0, 85, 170, 255] respectively) for testing. In summary, we employ 69,120 photo-realistic images for model training (consisting of 16 weather conditions $\times$ 30 car locations $\times$ 16 camera poses $\times$ 9 colors) and 49152 for testing (consisting of 16 weather conditions $\times$ 3 car locations $\times$ 16 camera poses $\times$ 64 colors)
@@
For texture generation and evaluation in physically-based simulation settings: For the generation of adversarial camouflage, we choose the same 16 weather combinations as NRP training. For evaluation in physically-based simulation, we produce two datasets: one with the same weather combinations as the training set and the other featuring 16 novel weather conditions that are not seen during the texture generation, including diverse scenarios like MidRainyNight and DustStorm. We use the former dataset for robust evaluation and the latter for transferability evaluation. Within each weather scenario, we randomly choose 20 locations for texture generation and 4 distinct locations for each test dataset. For each car transformation, imagery is captured at every 45-degree increment in azimuth angle, spanning four distinct altitude angles (0.0°, 22.5°, 45.0°, 67.5°) and four varying distances (5m, 10m, 15m, 20 m) for each. In summary, we employ 40,960 photo-realistic images for model training (consisting of 16 weather conditions $\times$ 20 car locations $\times$ 128 camera poses $\times$ 1 colors) and 7593 for each test dataset (consisting of 16 weather conditions $\times$ 4 car locations $\times$ 128 camera poses $\times$ 1 colors and obliterating the images of the car entirely obscured by the obstacle).
@@
For evaluation in the real world settings: We print our texture and adversarial texture from DAS (Wang et al., 2021), FCA (Wang et al., 2022), DTA (Suryanto et al., 2022) as well as ACTIVE (Suryanto et al., 2023) and stick them on official 1:12 Audi E-Tron vehicle models. For the adversarial texture generated by the UV-map-based method (DAS, FCA, and RAUCA), we print it and paste it on the vehicle directly according to its corresponding position since it is in the form of a UV map. For the world-align-based method (DTN and ACTIVE), since it generates a 2D square texture pattern, it cannot be directly pasted like the UV-map-based camouflage. Hence, we adopt the implementation from previous methods: repeating and enlarging the pattern, printing it, and applying it to the entire vehicle. In addition to this, we also prepare an unpainted vehicle. We shoot these models in the real world with the Xiaomi 14 for comparison. To demonstrate the robustness of our camouflage in physical multi-light conditions, we capture photos at different times to change the light condition. We take 2592 pictures of six cars with poses of 3 distances (5cm, 10cm, 15cm), 8 azimuths per 45°, and 3 altitude angles (0°, 30°, 60°), as well as at two locations (an asphalt road and an open square) at noon, in the afternoon, and at night . Moreover, we evaluate textures’ effectiveness in the presence of fog in the real world. We use a smoke machine to simulate the foggy environment on the imitated road on the desktop. We capture pictures of six cars at every 45-degree increment in azimuth angle, 3 altitude angles (0°, 30°, 60°). We take five consecutive pictures for the same camera angle with each car to reduce the unfairness caused by the randomness of the fog. Finally, we obtained 720 pictures for comparison.